# The swinging counterweight trebuchet Experiments on inner movement and ranges

### v1

@@ 2.2 Results of angular measurements
for three loads: A heavy metallic petanque ball to the left, a light tennis ball in the middle and a blank shot with no projectile to the right. The sensors were set to sample at a rate of 100Hz and to use a pre-trigger buffer such that data are collected from one second before $\theta_{e}$ passes $20^{0}$ and to stop two seconds later. The three seconds of acquisition time delivers 300 samples from well before the trebuchet is fired and until well after the projectile has left the pouch. All angles start at zero and broken lines show final values at rest for beam and counterweight. Oscillatory motion is seen in each case. It is strongest in the blank shot to the right, a little damped by the light projectile, and quite significantly suppressed by the heavy, which carries a large fraction of the available mechanical energy after release. The oscillations with smaller amplitude have larger frequency as expected for pendulum motion.
@@ 2.5 Motion of projectile and estimates of ranges, efficiencies and merits
Note that the path is not perpendicular to the sling because the spigot moves to the left. Figure 8b shows calculated vacuum ranges $R_{v}$ plotted as a function of horizontal release position. There is a clear maximum of $\simeq 62$m near $x=0.6$m. Spigot angles $\alpha$ are shown in figure 8c where it is seen that a setting close to $32^{0}$ leads to maximum range.
Similar results for the heavy projectile in figure 5 is shown in figure 9. The maximum range in vacuum is reduced to $\simeq 36$m and is found near $x=0.75$m for $\alpha\simeq 50^{0}$. Note that $\theta<\pi$ at the critical configuration while $\theta>\pi$ in figure 8a.
@@ 3.1 Varying projectile mass
The target is found at a negative value of $x$, so a positive release position subtracts from the range. The variable $y$ in figure 11b is the height at start measured from the base of the trestle, and it is added almost unchanged to the range for a target at ground level. The initial climb of the projectile $A$ in figure 11c is a little below 450 but approaches this value at large $m$. The speed at release $\textbf{v}|$ v |, which is the initial speed in the ballistic projectile path, is shown in figure 11d. It decreases with $m$ and reflects the same trend for the range $R_{v}$. The acceleration a projected on the velocity at release $a=\textbf{a}\cdot\textbf{v}/v$ is shown in figure 11e. It is quite large and positive for light projectiles so these are still gaining speed at release for longest range. The range maximizes at an initial climb $A$ of $45^{0}$ for constant speed, but since speed is increasing for decreasing $A$, a smaller $A$ is advantageous and figure 11c shows that $A\simeq 35^{0}$ is best for the lightest projectiles.
@@
The counterweight angle $\psi$ is small and negative at release, figure 12b, but during a shot it starts from zero and is first positive before it becomes negative. The angle between beam and counterweight, $\theta-\psi$, is shown in figure 12c. The configuration is stretched out when $\theta-\psi=\pi$, and this happens near $m=1$kg or close to maximum merit and an initial climb of $A=45^{0}$. The sling angle $\phi$ shown in figure 12d is near $45\pi/$ 4 at release, or $\pi/4=45^{0}$ over the horizon as seen also in figure 8 and 9.
@@ 3.4 Mechanical energy
The motion of sling (and projectile) is measured until the time when the pin P in figure 3a is stopped by the beam, but the projectile remains in the pouch a little longer. The motion of sling and projectile was calculated in this short interval of time by treating the system as a pendulum in an accelerated coordinate system that follows the spigot. The initial conditions for the motion was determined by the sling motion leading up to the pin stop, and it was followed until the direction of the projectile velocity dips below $-45^{0}$ relative to the horizon, such that a point just in front of the engine is hit. The extrapolation of the motion is short, but important, because mechanical energy is exchanged also here, and it allows all kinetic and potential energies in (6) or (7) to be calculated at all times.
@@ 3.7 Field measurements of range
Semi-empirical 0.87 717 37.5$\pm$0.1 35.1$\pm$0.5 50.00
predictions 0.87 1011 29.0$\pm$0.2 27.7$\pm$0.5 58.50
Field 0.87 700 34.4 $\pm$1.5 36.20
measurements 0.87 1011 24.9$\pm$1.5 32.50
0.93 700 37.5$\pm$1.5 19.70

### v2

@@ 2.2 Results of angular measurements
for three loads: A heavy metallic petanque ball to the left, a light tennis ball in the middle and a blank shot with no projectile to the right. The sensors were set to sample at a rate of 100Hz and to use a pre-trigger buffer such that data are collected from one second before $\theta_{e}$ passes $20^{\circ}$ and to stop two seconds later. The three seconds of acquisition time delivers 300 samples from well before the trebuchet is fired and until well after the projectile has left the pouch. All angles start at zero and broken lines show final values at rest for beam and counterweight. Oscillatory motion is seen in each case. It is strongest in the blank shot to the right, a little damped by the light projectile, and quite significantly suppressed by the heavy, which carries a large fraction of the available mechanical energy after release. The oscillations with smaller amplitude have larger frequency as expected for pendulum motion.
@@ 2.5 Motion of projectile and estimates of ranges, efficiencies and merits
Note that the path is not perpendicular to the sling because the spigot moves to the left. Figure 8b shows calculated vacuum ranges $R_{v}$ plotted as a function of horizontal release position. There is a clear maximum of $\simeq 62$m near $x=0.6$m. Spigot angles $\alpha$ are shown in figure 8c where it is seen that a setting close to $32^{\circ}$ leads to maximum range.
Similar results for the heavy projectile in figure 5 is shown in figure 9. The maximum range in vacuum is reduced to $\simeq 36$m and is found near $x=0.75$m for $\alpha\simeq 50^{\circ}$. Note that $\theta<\pi$ at the critical configuration while $\theta>\pi$ in figure 8a.
@@ 3.1 Varying projectile mass
The target is found at a negative value of $x$, so a positive release position subtracts from the range. The variable $y$ in figure 11b is the height at start measured from the base of the trestle, and it is added almost unchanged to the range for a target at ground level. The initial climb of the projectile $A$ in figure 11c is a little below 45 but approaches this value at large $m$. The speed at release $\textbf{v}|$ v |, which is the initial speed in the ballistic projectile path, is shown in figure 11d. It decreases with $m$ and reflects the same trend for the range $R_{v}$. The acceleration a projected on the velocity at release $a=\textbf{a}\cdot\textbf{v}/v$ is shown in figure 11e. It is quite large and positive for light projectiles so these are still gaining speed at release for longest range. The range maximizes at an initial climb $A$ of $45^{\circ}$ for constant speed, but since speed is increasing for decreasing $A$, a smaller $A$ is advantageous and figure 11c shows that $A\simeq 35^{\circ}$ is best for the lightest projectiles.
@@
The counterweight angle $\psi$ is small and negative at release, figure 12b, but during a shot it starts from zero and is first positive before it becomes negative. The angle between beam and counterweight, $\theta-\psi$, is shown in figure 12c. The configuration is stretched out when $\theta-\psi=\pi$, and this happens near $m=1$kg or close to maximum merit and an initial climb of $A=45^{\circ}$. The sling angle $\phi$ shown in figure 12d is near $45\pi/$ 4 at release, or $\pi/4=45^{\circ}$ over the horizon as seen also in figure 8 and 9.
@@ 3.4 Mechanical energy
The motion of sling (and projectile) is measured until the time when the pin P in figure 3a is stopped by the beam, but the projectile remains in the pouch a little longer. The motion of sling and projectile was calculated in this short interval of time by treating the system as a pendulum in an accelerated coordinate system that follows the spigot. The initial conditions for the motion was determined by the sling motion leading up to the pin stop, and it was followed until the direction of the projectile velocity dips below $-45^{\circ}$ relative to the horizon, such that a point just in front of the engine is hit. The extrapolation of the motion is short, but important, because mechanical energy is exchanged also here, and it allows all kinetic and potential energies in (6) or (7) to be calculated at all times.
@@ 3.7 Field measurements of range
Semi-empirical 0.87 717 37.5$\pm$0.1 35.1$\pm$0.5 50.0
predictions 0.87 1011 29.0$\pm$0.2 27.7$\pm$0.5 58.5
Field 0.87 700 34.4 $\pm$1.5 36.2
measurements 0.87 1011 24.9$\pm$1.5 32.5
0.93 700 37.5$\pm$1.5 19.7